OpenStudy (anonymous):

Solve: x - 6 = -12 A) -18 Eliminate B) -6 C) 2 D) 72

OpenStudy (jiteshmeghwal9):

add 6 both sides :)

OpenStudy (anonymous):

positive six right

OpenStudy (anonymous):

B. x-6=-12 add 6 to both sides. x-6+6 = -12+6 -12+6 = -6 x=-6

OpenStudy (jiteshmeghwal9):

right

OpenStudy (anonymous):

yea its gona come out t negative six because you are going to use the - sign because it is larger

OpenStudy (jiteshmeghwal9):

@TroubleMakerBrit don't post direct answer, it's against the code of conduct :)

OpenStudy (anonymous):

thanks though @TroubleMakerBrit and @jiteshmeghwal9 :}}

OpenStudy (jiteshmeghwal9):

Let the ask er to participate in the question :)

OpenStudy (jiteshmeghwal9):

yw :)